OpenStudy (anonymous):

find the inverse of the function f(x)=x^2+x-7

OpenStudy (tkhunny):

What's keeping you from the usual "Swap and Solve" idea? Note: It won't be pretty. The inverse will NOT be a function.

OpenStudy (anonymous):

why?

OpenStudy (tkhunny):

x = y^{2} + y - 7 -- That was the Swap Now, solve for y. Complete the square, Quadratic formula, whatever comes to mind.

OpenStudy (anonymous):

y=x+7/y+1

OpenStudy (anonymous):

??

OpenStudy (tkhunny):

That is not solving for y. \(x = y^{2} + y - 7\) \(y^{2} + y - (7+x) = 0\) \(y = \dfrac{-1\pm\sqrt{1 + 4(7+x)}}{2}\) No, THAT is solving for y. just like any other quadratic equation.